# Structured and Balanced Multi-component and Multi-layer Networks

### v1

@@ 2.4 MMNNs versus FCNNs
$(a_{i,j})=\begin{bmatrix}0.3&0.2\\ 0.2&0.3\end{bmatrix},\qquad(b_{i})=\begin{bmatrix}2\pi\\ 4\pi\end{bmatrix},\qquad(c_{i,j})=\begin{bmatrix}2\pi&4\pi\\ 8\pi&4\pi\end{bmatrix},\quad\textnormal{and}\quad(d_{i,j})=\begin{bmatrix}4\pi&6\pi\\ 8\pi&6\pi\end{bmatrix}.$
@@
Figure 3: $f_{1}$.
Figure 4: $f_{2}$.
We select large network sizes (see Table 2) to ensure that all networks possess sufficient parameters to learn the target functions. [^2] FCNNs perform poorly if the network size is small. For a fair comparison, we choose relatively large network sizes for FCNNs and MMNNs, where both perform reasonably well. For training the 1D function, we sample a total of 1000 data points on a uniform grid within $[-1,1]$, using a mini-batch size of 100 and a learning rate of $0.001\times 0.9^{\lfloor k/400\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},20000$. For training the 2D function, we sample a total of $600^{2}$ data points on a uniform grid within $[-1,1]^{2}$, using a mini-batch size of 1000 and a learning rate of $0.001\times 0.9^{\lfloor k/16\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},800$.
@@
$f_{1}$ MMNN1 (S1) (388, 18, 6) 35399 / 73035 $2.49\times 10^{-6}$ $9.93\times 10^{-3}$ 23.3s / 1000 epochs
@@
$f_{1}$ MMNN1 (S2) (388, 18, 6) 73035 / 73035 $2.05\times 10^{-6}$ $1.88\times 10^{-2}$ 27.4s / 1000 epochs
@@
$f_{2}$ MMNN2 (S1) (789, 36, 12) 313630 / 637120 $4.61\times 10^{-6}$ $1.55\times 10^{-2}$ 30.3s / 10 epochs
@@
Figure 5: First row: logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$. Second row: differences between learned networks and $f_{1}$. Second row: differences between learned networks and $f_{2}$.
@@ 3 Multi-component and multi-layer decomposition
Although a one-hidden-layer neural network is a low-pass filter that can not represent and learn high-frequency features effectively [38], we use mathematical construction to show that MMNNs, which are composed of one-hidden-layer neural networks, can overcome this difficulty by decomposition of the complexity through components and/or depth. We emphasize that the decomposition is highly non-unique. Our construction is “man-made” which can be different from the one by computer through an optimization (learning) process. Our discussion begins with one-dimensional construction in Section 3.1 and later extends to higher dimensions in Section 3.2.
@@ 3.1 One dimensional construction
We begin with a two-component decomposition in 1D as both an illustration and an example in Section 3.1.1. Later in Section 3.1.2, we introduce the general multi-component decomposition. Finally in Section 3.1.3, we use concrete examples for demonstration.
@@ 3.1.1 Two-component decomposition
We show a simple “divide and conquer” strategy for a target function (example) $f(x)=\cos(2n\pi x)$, a high frequency Fourier mode when $n$ is large. Define
${\bm{f}}_{1}:[-1,1]\mapsto[-1,1]^{2},\qquad{\bm{f}}_{1}=\begin{bmatrix}f_{1,1}\\ f_{1,2}\end{bmatrix},$
$f_{1,1}(x)={\mathtt{ReLU}}(2x)-1=\begin{cases}-1&\textnormal{for}\ x\in[-1,0),\\ 2x-1&\textnormal{for}\ x\in[0,1],\end{cases}$
$f_{1,2}(x)=-{\mathtt{ReLU}}(-2x)+1=\begin{cases}2x+1&\textnormal{for}\ x\in[-1,0),\\ 1&\textnormal{for}\ x\in[0,1],\end{cases}$
@@
$\begin{split}f(x)&=\cos\Big{(}n\pi\cdot{\mathtt{ReLU}}(2x)\Big{)}+\cos\Big{(}-n\pi\cdot{\mathtt{ReLU}}(-2x)\Big{)}\\ &=\cos\Big{(}n\pi\big{(}f_{1,1}(x)+1\big{)}\Big{)}+\cos\Big{(}n\pi\big{(}f_{1,2}(x)-1\big{)}\Big{)}=f_{2}\circ{\bm{f}}_{1}(x)\end{split}$
@@
${\bm{f}}_{1}:[-1,1]\mapsto[-1,1]^{2},\qquad{\bm{f}}_{1}=\begin{bmatrix}f_{1,1}\\ f_{1,2}\end{bmatrix},$
where the component functions $f_{1,1}$ and $f_{1,2}$ are defined by
$f_{1,1}(x)={\mathtt{ReLU}}(2x)-1=\begin{cases}-1&\textnormal{for}\ x\in[-1,0),\\ 2x-1&\textnormal{for}\ x\in[0,1],\end{cases}$
and
$f_{1,2}(x)=-{\mathtt{ReLU}}(-2x)+1=\begin{cases}2x+1&\textnormal{for}\ x\in[-1,0),\\ 1&\textnormal{for}\ x\in[0,1].\end{cases}$
Moreover,
@@
Hence, for any $x\in[-1,1]$, we achieve the following reconstruction of $f(x)$:
@@ 3.1.2 General multi-component decomposition
Now we propose a general multi-component adaptive decomposition, a “divide and conquer” strategy, that can distribute the complexity of a target function evenly to multiple components.
@@
${\mathcal{L}}_{i}(a_{i})=x_{i-1}\quad\textnormal{and}\quad{\mathcal{L}}_{i}(b_{i})=x_{i}.$ (2)
@@
$f_{i}=f\circ{\mathcal{L}}_{i}:[a_{i},b_{i}]\to{\mathbb{R}}.$ (3)
@@
$\psi_{i}(x)=s_{i}\cdot{\mathtt{ReLU}}\left({x-x_{i-1}}\right)-s_{i}\cdot{\mathtt{ReLU}}\left({x-x_{i}}\right)+a_{i}.$ (4)
@@
$\begin{split}\sum_{i=1}^{n}f_{i}\circ\psi_{i}(x)&=\sum_{i=1}^{n}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)=\sum_{i=1}^{j-1}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)+f\circ{\mathcal{L}}_{j}\circ\psi_{j}(x)+\sum_{i=j+1}^{n}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)\\ &=\sum_{i=1}^{j-1}f\circ{\mathcal{L}}_{i}(b_{i})+f\circ{\mathcal{L}}_{j}\circ{\mathcal{L}}_{j}^{-1}(x)+\sum_{i=j+1}^{n}f\circ{\mathcal{L}}_{i}(a_{i})\\ &=\sum_{i=1}^{j-1}f(x_{i})+f(x)+\sum_{i=j+1}^{n}f(x_{i-1})=f(x)+\underbrace{\sum_{i=1}^{n-1}f(x_{i})}_{\textnormal{constant}}.\end{split}$
@@
Since $j$ is arbitrary, the above equation holds for all $x=\cup_{j=1}^{n}[x_{j-1},x_{j}]=[x_{0},x_{n}]$. ∎

### v2

@@ 3.1 Oscillatory function approximation
$(a_{i,j})=\begin{bmatrix}0.3&0.2\\ 0.2&0.3\end{bmatrix},\qquad(b_{i})=\begin{bmatrix}2\pi\\ 4\pi\end{bmatrix},\qquad(c_{i,j})=\begin{bmatrix}2\pi&4\pi\\ 8\pi&4\pi\end{bmatrix},\quad(d_{i,j})=\begin{bmatrix}4\pi&6\pi\\ 8\pi&6\pi\end{bmatrix}.$
@@
Figure 3: Plot of $f_{1}$.
Figure 4: Plots of $f_{2}$.
Large network sizes (see Table 2) are selected to ensure that all networks possess sufficient parameters to learn the target functions. [^2] FCNNs perform poorly if the network size is small. For a fair comparison, we choose relatively large network sizes for FCNNs and MMNNs, where both perform reasonably well. For training the one-dimensional function, we sample a total of 1000 data points on a uniform grid within $[-1,1]$, using a mini-batch size of 100 and a learning rate of $0.001\times 0.9^{\lfloor k/400\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},20000$. For training the two-dimensional function, we sample a total of $600^{2}$ data points on a uniform grid within $[-1,1]^{2}$, using a mini-batch size of 1000 and a learning rate of $0.001\times 0.9^{\lfloor k/16\rfloor}$ for epochs $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},800$. The Adam optimizer is employed for both functions.
@@
$f_{1}$ MMNN1 (S1) (388, 18, 6) 35399 / 73035 $2.49\times 10^{-6}$ $\bm{9.93\times 10^{-3}}$ 23.3s / 1000 epochs
@@
$f_{1}$ MMNN1 (S2) (388, 18, 6) 73035 / 73035 $\bm{2.05\times 10^{-6}}$ $1.88\times 10^{-2}$ 27.4s / 1000 epochs
@@
$f_{2}$ MMNN2 (S1) (789, 36, 12) 313630 / 637120 $\bm{4.61\times 10^{-6}}$ $\bm{1.55\times 10^{-2}}$ 30.3s / 10 epochs
@@
Figure 5: First row: logarithm of “test-error-aver” vs. epoch, where “test-error-aver” for epoch $k$ is calculated by averaging the errors in epochs $\max\{1,k-100\}$ to $\min\{k+100,\#\text{epochs}\}$. All errors shown on the $y$-axis are in base-10 logarithmic scale. Second row: differences between learned networks and $f_{1}$. Third row: differences between learned networks and $f_{2}$.
@@
3.2 Solving partial differential equations
Next, we compare the performance of MMNNs and FCNNs for solving partial differential equations (PDEs). We consider a classical example: the two-dimensional Poisson equation with zero Dirichlet boundary conditions:
$-\Delta u(x,y)=f(x,y),\quad(x,y)\in(-1,1)^{2},\quad u|_{\partial\Omega}=0,$
where the source term is given by
$f(x,y)=-112\pi^{2}\sin(7\pi x)\sin(8\pi y)-108\pi^{2}\sin(6\pi x)\sin(9\pi y).$
It is easy to verify that the exact solution is
$u(x,y)=\sin(7\pi x)\sin(8\pi y)+\sin(6\pi x)\sin(9\pi y).$
We approximate the solution $u(x,y)$ by a neural network $u_{\bm{\theta}}(x,y)$, where ${\bm{\theta}}$ denotes the network parameters and solve the Poisson equation using the Physics-Informed Neural Network (PINN) [27] formulation. The PDE residual is defined as $\mathcal{R}(x,y)=-\Delta u_{\bm{\theta}}(x,y)-f(x,y)$, where $\Delta u_{\bm{\theta}}=\frac{\partial^{2}u_{\bm{\theta}}}{\partial x^{2}}+\frac{\partial^{2}u_{\bm{\theta}}}{\partial y^{2}}$, and the neural network is trained to minimize the loss function $\mathcal{L}_{\text{total}}=\lambda_{\text{PDE}}\mathcal{L}_{\text{PDE}}+\lambda_{\text{BC}}\mathcal{L}_{\text{BC}}$, combining the PDE loss $\mathcal{L}_{\text{PDE}}=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\mathcal{R}(x_{i},y_{i})^{2}$ for collocation points $(x_{i},y_{i})\in(-1,1)^{2}$ and the boundary loss $\mathcal{L}_{\text{BC}}=\frac{1}{N_{b}}\sum_{j=1}^{N_{b}}u_{\bm{\theta}}(x_{j},y_{j})^{2}$ for boundary points $(x_{j},y_{j})\in\partial\Omega$.
To compare FCNNs and MMNNs in terms of representation accuracy, we avoid soft boundary condition enforcement by including the penalty term $\lambda_{\text{BC}}\mathcal{L}_{\text{BC}}$ in the loss, which requires careful tuning of $\lambda_{\text{BC}}$. We adopt a hard-constraint formulation that guarantees the boundary condition is satisfied exactly. Specifically, we define the network output as
$u_{\bm{\theta}}(x,y)=h_{\bm{\theta}}(x,y)\cdot\cos\left(\frac{\pi x}{2}\right)\cos\left(\frac{\pi y}{2}\right),$
where $h_{\bm{\theta}}(x,y)$ is modeled by either an FCNN or an MMNN. This construction ensures that $u_{\bm{\theta}}(x,y)=0$ on the boundary $\partial\Omega$ for all values of $\bm{\theta}$.
We select an MMNN of size (301, 16, 6), another MMNN of size (503, 20, 6), and an FCNN of size (100, –, 6), denoted as MMNN1, MMNN2, and FCNN, respectively, for simplicity. Since $\mathtt{ReLU}$ is not differentiable, we use the sine function as the activation function. We sample $100^{2}$ data points for $f(x,y)$ on a uniform grid in $(-1,1)^{2}$, using a mini-batch size of 2000 and setting $\lambda_{\text{PDE}}=0.001$. We adopt the Adam optimizer, with the learning rate set to $\lfloor k/200\rfloor/800$ for $k<16000$, and to $0.001\times 0.9^{\lfloor(k-16000)/1600\rfloor}$ for $k\geq 16000$, where $k=1,2,\mathinner{\mkern-0.1mu\cdotp\mkern-0.3mu\cdotp\mkern-0.3mu\cdotp\mkern-0.1mu},160000$ denotes the training epoch. We note that for $k<16000$, we use an increasing learning rate to facilitate warm-up (see, e.g., [1, 15]), which enhances training performance.
Our experiments reveal that initial parameters significantly impact training, particularly for FCNNs. To ensure experimental reliability, we repeated the experiments with 16 different seeds. As demonstrated in Figures 6, 7, and Table 3, both MMNNs surpass the FCNN in solving PDEs with PINNs, even with comparable depth and total (or training) parameters. We note that using the same seed may yield different outcomes across various code environments. Our tests consistently show that MMNNs always succeed, whereas FCNNs fail with high probability (increasing FCNN width may improve the likelihood of success).
(a) Training error.
(b) Test error (MSE).
(c) Test error (MAX).
Figure 6: Average errors across 16 seeds versus epoch. The training error corresponds to the PDE loss, while the test error (MSE or MAX) quantifies the difference between the learned network and the true solution. All errors shown on the $y$-axis are in base-10 logarithmic scale.
(a) Truth.
(b) MMNN1.
(c) MMNN2.
(d) FCNN.
Figure 7: Comparison of three networks: (a) true solution; (b, c, d) differences between the true solution and predictions from learned networks. For each network, we select the best trained model from 16 seeds.
Table 3: Comparison of test errors.
#parameters (trained / all) 24462 / 50950 50904 / 105228 50901 / 50901
MMNN1 of size (301, 16, 6) MMNN2 of size (503, 20, 6) FCNN of size (100, –, 6)
seed for randomness MSE MAX MSE MAX MSE MAX
00 $2.60\times 10^{-6}$ $6.22\times 10^{-3}$ $1.77\times 10^{-8}$ $3.38\times 10^{-3}$ $4.95\times 10^{-1}$ $1.94\times 10^{0}$
$1$ $2.18\times 10^{-7}$ $2.08\times 10^{-3}$ $9.32\times 10^{-9}$ $6.76\times 10^{-4}$ $\bm{7.95\times 10^{-7}}$ $3.35\times 10^{-3}$
$2$ $1.18\times 10^{-6}$ $4.18\times 10^{-3}$ $2.38\times 10^{-8}$ $1.55\times 10^{-3}$ $5.01\times 10^{-1}$ $1.94\times 10^{0}$
$3$ $3.19\times 10^{-6}$ $5.57\times 10^{-3}$ $6.68\times 10^{-9}$ $1.19\times 10^{-3}$ $5.03\times 10^{-1}$ $1.94\times 10^{0}$
$4$ $5.34\times 10^{-7}$ $3.12\times 10^{-3}$ $2.35\times 10^{-8}$ $2.89\times 10^{-3}$ $4.99\times 10^{-1}$ $1.94\times 10^{0}$
$5$ $1.76\times 10^{-6}$ $5.15\times 10^{-3}$ $1.78\times 10^{-8}$ $2.51\times 10^{-3}$ $4.96\times 10^{-1}$ $1.94\times 10^{0}$
$6$ $4.92\times 10^{-7}$ $2.61\times 10^{-3}$ $9.68\times 10^{-9}$ $4.65\times 10^{-3}$ $5.04\times 10^{-1}$ $1.94\times 10^{0}$
$7$ $1.97\times 10^{-6}$ $5.34\times 10^{-3}$ $3.87\times 10^{-8}$ $3.37\times 10^{-3}$ $5.01\times 10^{-1}$ $1.94\times 10^{0}$
$8$ $4.98\times 10^{-7}$ $2.44\times 10^{-3}$ $4.04\times 10^{-8}$ $9.83\times 10^{-4}$ $8.20\times 10^{-7}$ $\bm{3.22\times 10^{-3}}$
$9$ $2.57\times 10^{-6}$ $4.10\times 10^{-3}$ $\bm{1.69\times 10^{-9}}$ $\bm{6.66\times 10^{-4}}$ $5.00\times 10^{-1}$ $1.94\times 10^{0}$
$10$ $2.47\times 10^{-6}$ $6.15\times 10^{-3}$ $2.72\times 10^{-8}$ $3.88\times 10^{-3}$ $5.02\times 10^{-1}$ $1.94\times 10^{0}$
$11$ $1.03\times 10^{-6}$ $3.86\times 10^{-3}$ $2.02\times 10^{-8}$ $1.31\times 10^{-3}$ $5.03\times 10^{-1}$ $1.94\times 10^{0}$
$12$ $1.61\times 10^{-6}$ $5.13\times 10^{-3}$ $4.38\times 10^{-9}$ $1.16\times 10^{-3}$ $1.60\times 10^{-5}$ $1.12\times 10^{-2}$
$13$ $\bm{9.35\times 10^{-8}}$ $\bm{1.52\times 10^{-3}}$ $3.85\times 10^{-8}$ $1.07\times 10^{-2}$ $2.86\times 10^{-2}$ $4.39\times 10^{-1}$
$14$ $7.42\times 10^{-7}$ $3.41\times 10^{-3}$ $1.51\times 10^{-8}$ $1.77\times 10^{-3}$ $5.01\times 10^{-1}$ $1.95\times 10^{0}$
$15$ $6.58\times 10^{-6}$ $6.25\times 10^{-3}$ $2.74\times 10^{-8}$ $1.23\times 10^{-3}$ $6.85\times 10^{-6}$ $7.97\times 10^{-3}$
@@ 4 Multi-component and multi-layer decomposition
It has been shown in [41] that a one-hidden-layer neural network acts as a low-pass filter and cannot effectively represent or learn high-frequency features. Using mathematical construction, we demonstrate that MMNNs, which are composed of one-hidden-layer neural networks, can overcome this difficulty by decomposing complexity through their components and/or depth. We emphasize that the decomposition is highly non-unique. Our construction is “man-made” which can be different from the one by computer through an optimization (learning) process. Our discussion begins with one-dimensional construction in Section 4.1 and later extends to higher dimensions in Section 4.2.
@@ 4.1 One dimensional construction
We begin with a two-component decomposition in the one-dimensional as both an illustration and an example in Section 4.1.1. Later in Section 4.1.2, we introduce the general multi-component decomposition. Finally in Section 4.1.3, we use concrete examples for demonstration.
@@ 4.1.1 Two-component decomposition
We demonstrate a simple “divide and conquer” strategy for an example function $f(x)=\cos(2n\pi x)$, a high frequency Fourier mode when $n$ is large. Define
and ${\bm{f}}_{1}=\begin{bmatrix}f_{1,1}\\ f_{1,2}\end{bmatrix}:[-1,1]\mapsto[-1,1]^{2}$, where the components $f_{1,1}$ and $f_{1,2}$ are given by
(2) $f_{1,1}(x)={\mathtt{ReLU}}(2x)-1=\begin{cases}-1&\textnormal{for}\ x\in[-1,0),\\ 2x-1&\textnormal{for}\ x\in[0,1],\end{cases}$
@@
(3) $f_{1,2}(x)=-{\mathtt{ReLU}}(-2x)+1=\begin{cases}2x+1&\textnormal{for}\ x\in[-1,0),\\ 1&\textnormal{for}\ x\in[0,1].\end{cases}$
@@
$\begin{split}f(x)&=\cos\Big{(}n\pi\cdot{\mathtt{ReLU}}(2x)\Big{)}+\cos\Big{(}-n\pi\cdot{\mathtt{ReLU}}(-2x)\Big{)}\\ &=\cos\Big{(}n\pi\big{(}f_{1,1}(x)+1\big{)}\Big{)}+\cos\Big{(}n\pi\big{(}f_{1,2}(x)-1\big{)}\Big{)}=f_{2}\circ{\bm{f}}_{1}(x).\end{split}$
@@
and ${\bm{f}}_{1}=\begin{bmatrix}f_{1,1}\\ f_{1,2}\end{bmatrix}:[-1,1]\mapsto[-1,1]^{2}$, where $f_{1,1}$ and $f_{1,2}$ are given in (2) and (3). Hence, for any $x\in[-1,1]$, we achieve the following reconstruction of $f(x)$:
@@ 4.1.2 General multi-component decomposition
Now we extend to a general multi-component adaptive decomposition, a “divide and conquer” strategy, that can distribute the complexity of a target function evenly to multiple components.
@@
(4) ${\mathcal{L}}_{i}(a_{i})=x_{i-1}\quad\textnormal{and}\quad{\mathcal{L}}_{i}(b_{i})=x_{i}.$
@@
(5) $f_{i}=f\circ{\mathcal{L}}_{i}:[a_{i},b_{i}]\to{\mathbb{R}}.$
@@
(6) $\psi_{i}(x)=s_{i}\cdot{\mathtt{ReLU}}\left({x-x_{i-1}}\right)-s_{i}\cdot{\mathtt{ReLU}}\left({x-x_{i}}\right)+a_{i}.$
@@
$\begin{split}\sum_{i=1}^{n}f_{i}\circ\psi_{i}(x)&=\sum_{i=1}^{n}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)\\ &=\sum_{i=1}^{j-1}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)+f\circ{\mathcal{L}}_{j}\circ\psi_{j}(x)+\sum_{i=j+1}^{n}f\circ{\mathcal{L}}_{i}\circ\psi_{i}(x)\\ &=\sum_{i=1}^{j-1}f\circ{\mathcal{L}}_{i}(b_{i})+f\circ{\mathcal{L}}_{j}\circ{\mathcal{L}}_{j}^{-1}(x)+\sum_{i=j+1}^{n}f\circ{\mathcal{L}}_{i}(a_{i})\\ &=\sum_{i=1}^{j-1}f(x_{i})+f(x)+\sum_{i=j+1}^{n}f(x_{i-1})=f(x)+\underbrace{\sum_{i=1}^{n-1}f(x_{i})}_{\textnormal{constant}}.\end{split}$
@@
Since $j$ is arbitrary, the above equation holds for all $x=\cup_{j=1}^{n}[x_{j-1},x_{j}]=[x_{0},x_{n}]$.